# A Survey on Computational Solutions for Reconstructing Complete Objects by Reassembling Their Fractured Parts

### v1

@@ 8.2 Software Packages
Few early approaches to fractured object reassembly released code. A major factor is that in the early days we did not have established benchmark datasets for experimental evaluation. For example, there are rich benchmark datasets in computer vision and people deliver open-source codes that execute on benchmark datasets. With the introduction of large-scale benchmark datasets, we see the trend to release codes. For example, many papers that evaluate using Breaking Bad have open-source implementations, including Jigsaw [61], SE(3) equivariance [142], DiffAssemble [153], and PuzzleFusion++ [141].
@@ 9 Future Directions
On the representation and algorithm side, there are open questions on how to learn shape priors of diverse complete objects. When targeting general object categories that have large topological and geometrical variations, implicit shape representations, which can encode shapes with varying topology, become the only choice. However, it is an open problem to properly define correspondences between fragments, which are normally represented as point clouds.
@@
In addition, having sufficient training data is critical to the success of developing deep neural networks. However, this becomes challenging in the case of the fragment reassembly problem. First of all, the number of complete objects that have fragments remain sparse. In addition, developing software and hardware systems to scan these fragments is very costly. There are two interesting topics related to addressing this data issue.
The first topic is on how to boost the performance of reassembly using vision foundation models. Many sub-tasks in the reassembly pipeline, including fracture surface segmentation, establishing correspondences between pairs of fragments, predicting the underlying complete object from fragments, and matching fragments with a complete object, have counterparts as core geometry processing and 3D vision problems. Recently, we have seen that 2D foundation models, either through rendered images or lifted 2D features, have led to breakthrough results on these tasks. This brings interesting questions on how to adopt these approaches on the corresponding sub-tasks in reassembly. For example, a recent paper [300] has shown that vision-language models can effectively understand the material properties of objects from rendered images. Note that rendering images of fragments is fundamentally different from rendering images of complete objects for visual understanding. How to do so to reveal fracture regions is a very interesting research question.
@@
Once we have an end-to-end learnable fracture simulation pipeline, another important problem is to study how to combine fracture reassembly and fracture simulation in an end-to-end manner. This allows us to optimize the fracture simulation network, so that the fracture reassembly network generalizes optimally when it is trained from the fracture simulation output. This task goes beyond the standard notion of end-to-end training, as we have to develop losses on the output distribution of the fracture simulation network.
From a community-building perspective, it is important to create large-scale benchmark datasets for experimental evaluation. This will attract more people to work on this problem and, according to the trend in computer vision, will simulate more people to release open-source codes. We do observe a good initial trend from the Breaking Bad dataset [145].
@@ 10 Conclusions
In this survey, we have provided a thorough survey on the reassembly problem, which finds rich applications in cultural heritage restoration, bone fracture restoration, and object/scene modeling from parts and objects. We have covered early non-deep learning based approaches to more recent deep learning based approaches. Due to connections to core geometry processing problems in shape segmentation, shape matching, symmetry and primitive detection, and learning shape priors, we also discussed relevant approaches on these problems, emphasizing similarities and differences. We also covered current state-of-the-art methods for solving real problems in these application domains.
@@
Acknowledgement. Qixing Huang is supported by NSF Career IIS-2047677, NSF IIS-2413161, and Gifts from Adobe and Google. Xin Li is partly supported by NSF CBET-2115405 and NIH R15HD108765. Junfeng Jiang would like to acknowledge support from Jiangsu Province Key Research Program-Social Development-Clinical Frontier Technologies (BE2022718).

### v2

@@ 8.3 Open Source
Few early approaches to fractured object reassembly released code. A major factor is that in the early days, we did not have established benchmark datasets for experimental evaluation. For example, there are rich benchmark datasets in computer vision and researchers deliver open-source codes that execute on benchmark datasets. With the introduction of large-scale benchmark datasets, we see the trend to release codes. For example, many papers that evaluate using Breaking Bad have open-source implementations, including Jigsaw [62], SE(3) equivariance [146], PHFormer [154], DiffAssemble [157], PMTR [151], and PuzzleFusion++ [145]. We provide a comprehensive benchmark performance on these open-sourced methods in Table 4. For procedural methods such as ICP on the large-scale dataset, we encourage the readers to refer to the results presented in Neural Shape Mating [203].
@@ 9 Future Directions
On the representation and algorithm side, there are open questions on how to learn shape priors of diverse complete objects. When targeting general object categories that have large topological and geometrical variations, implicit shape representations, which can encode shapes with varying topology, become the only choice. However, establishing reliable correspondences between implicit representations of complete shapes and the underlying geometric structure of fragments remains an open research challenge.
@@
In addition, having sufficient training data is critical to the success of developing deep neural networks. However, this becomes challenging in the case of the fragment reassembly problem. First of all, the number of complete objects that have fragments remains sparse. In addition, developing software and hardware systems to scan these fragments is very costly. There are two interesting topics related to addressing this data issue.
The first topic explores boosting reassembly performance using vision foundation models. Many subtasks in the reassembly pipeline - including fracture surface segmentation, establishing correspondences between pairs of fragments, predicting the underlying complete object from fragments, and matching fragments with a complete object - have counterparts in core geometry processing and 3D vision problems. However, these tasks face unique challenges in the reassembly context. For correspondence establishment, current methods primarily rely on geometric matching using surface features or contours, which shows limited generalization ability across different object categories and fracture patterns. The assembly problem requires both geometric matching between fragments and semantic matching between fragments and templates - a dual objective that few existing approaches address. Additionally, while there has been significant progress in 3D shape generation using implicit representations and neural fields, predicting complete objects from fragments remains challenging. Most point cloud generation methods struggle with the diversity of possible completions, especially when handling rotated or partially observed objects. Recently, we have seen that 2D foundation models, either through rendered images or lifted 2D features, have led to breakthrough results on these tasks. 2D pretrained model, e.g. DINOv2 [309] and Stable Diffusion [310], have shown potential as depth estimation backbone. The extracted features have also shown success in finding correspondences [311, 312, 313] with a 30% improvement. This brings interesting questions on how to adopt these approaches in the corresponding subtasks in reassembly. For example, a recent paper [314] has shown that vision-language models can effectively understand the material properties of objects from rendered images. However, rendering fragments presents unique challenges compared to complete objects, as fracture regions need special consideration for effective visual understanding. Developing rendering techniques that can reveal and emphasize fracture characteristics while maintaining other geometric and material properties remains an open research question.
@@
Once we have an end-to-end learnable fracture simulation pipeline, another important problem is studying how to combine fracture reassembly and simulation in an end-to-end manner. This allows optimizing the simulation network so the reassembly network generalizes optimally when trained from simulation output. This task goes beyond standard end-to-end training, as we must develop losses on the output distribution of the simulation network.
From a community-building perspective, it is important to create large-scale benchmark datasets for experimental evaluation. This will attract more researchers to work on this problem and, according to the trend in computer vision, will stimulate more researchers to release open-source codes. We do observe a good initial trend from the Breaking Bad dataset [149].
@@ 10 Conclusions
In this survey, we have provided a thorough survey of the reassembly problem, which finds rich applications in cultural heritage restoration, bone fracture restoration, and object/scene modeling from parts and objects. We have covered early procedural approaches to more recent deep learning based approaches. Due to connections to core geometry processing problems in shape segmentation, shape matching, symmetry and primitive detection, and learning shape priors, we also discussed relevant approaches to these problems, emphasizing similarities and differences. We also covered current state-of-the-art methods for solving real problems in these application domains.
@@
Acknowledgement
Qixing Huang is supported by NSF Career IIS-2047677, NSF IIS-2413161, and Gifts from Adobe and Google. Xin Li is partly supported by NSF-IIS 1320959 and NSF CBET-2115405. Junfeng Jiang would like to acknowledge support from Jiangsu Province Key Research Program-Social Development-Clinical Frontier Technologies (BE2022718).